40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sat Sep 24 01:21:11 2011, kiwamu, Update, SUS, damping test 
    Reply  Mon Sep 26 14:21:07 2011, kiwamu, Update, SUS, ITMX ULSEN shows jumps ITMXspectra.pngITMX-sensors.png
       Reply  Mon Sep 26 15:54:46 2011, kiwamu, Update, SUS, ITMX ULSEN shows jumps 
          Reply  Mon Sep 26 16:42:08 2011, kiwamu, Update, SUS, ITMX ULSEN : fixed connection.pngITMXsensors.png
             Reply  Wed Sep 28 07:46:42 2011, steve, Update, SUS, Dsub connections at the back of 1X5 are fixed 
Message ID: 5544     Entry time: Mon Sep 26 14:21:07 2011     In reply to: 5534     Reply to this: 5546
Author: kiwamu 
Type: Update 
Category: SUS 
Subject: ITMX ULSEN shows jumps 

Quote from #5534

As a suspension test I am leaving all of the suspensions restored and damped with OSEMS but without oplevs

According to the spectra, all of the suspensions had been damped with the OSEMs. The peaks around 1Hz are reasonably suppressed.

However the spectra from ITMX showed a noise floor at very high level. This is because of strange jumps in the signal of the UL shadow sensor.

I will check some analog circuits for the UL sensor.

 


(ITMX shadow sensors)

Here is the spectra of the ITMX shadow sensors taken during the damping test (#5534)- -

 ITMXspectra.png 

The UL sensor shows a unacceptable amount of noise.

 

Additionally I checked the time series of the ITMX shadow sensors and found ONLY the UL sensor frequently showed strange jumps in data.

Here is an example of the time series showing a jump ONLY in the UL sensor.

ITMX-sensors.png

It is possible that the jumps are coming from some circuits, since the rest of the sensors (including the oplevs) don't detect the same jump.

ELOG V3.1.3-